The sum of three numbers is 98. The first number is 2/3 of the second, and the second is 5/8 of the third. What is the second number?
Answer by checkley71(8403)

x+y+z=98
x=2y/3
y=5z/8 or 5z=8y or z=8y/5
2y/3+y+8y/5=98
(5*2y+15y+3*8y)/15=98
(10y+15y+24y)/15=98 now cross multiply
49y=15*98
49y=1470
y=1470/49
y=30 the second number.
x=2*30/3=60/3=20 for the first number.
30+20+z=98
50+z=98
z=98-50
z=48 the third number.